# When do World Models Successfully Learn Dynamical Systems?

### v1

@@ A.1 Extraneous Proofs
A.2 Continuous Time Control Theory
In continuous time control theory, Definition 1 (i) is changed by demanding $y(t;x_{0})=y(t;x_{0})$ for $t\in[0,T]$ instead of only for $t\in\Gamma_{T}$. Note that the dynamical system (1) is a special case of [45] since the distinguishing control is assumed to be the trivial one. Let us emphasize that for nonlinear systems this generally diminishes otherwise existent observability properties, cf. [60, Section 5]. On the other hand, considering uncontrolled systems has the advantage of characterizing observability equivalently in terms of the injectivity of the so-called output map
$H\colon\mathcal{X}\to L^{2}(0,T;\mathbb{R}^{m}),\ \ H(x_{0}):=h(x(\cdot;x_{0}))=y(\cdot;x_{0})$ (23)
which assigns to a given initial state its corresponding output trajectory. As a consequence, observability of $(f,h)$ implies the existence of an inverse $H^{-1}\colon H(\mathcal{X})\to\mathcal{X}$. Composing this function with the flow of (1) moreover allows to define a reconstruction map
$G\colon H(\mathcal{X})\to L^{2}(0,\tilde{T};\mathbb{R}^{n}),\ \ G(y):=(S_{\tilde{T}}\circ H^{-1})(y)=x(\cdot;H^{-1}(y))$ (24)
@@ A.2 Continuous Time Control Theory
$H^{-1}y=Q(\tilde{T})^{-1}\int_{0}^{\tilde{T}}e^{A^{\top}t}h^{\top}y(t)\,\mathrm{d}t,\quad Gy=e^{A\cdot}Q(\tilde{T})^{-1}\int_{0}^{\tilde{T}}e^{A^{\top}t}h^{\top}y(t)\,\mathrm{d}t$ (25)
where the positive definiteness of the observability Gramian $Q(\tilde{T}):=\int_{0}^{\tilde{T}}e^{A^{\top}s}h^{\top}he^{As}\,\mathrm{d}s$ is ensured by the observability of the system, see [46, Section 6.3].
In [45], other notions of observability have been discussed. Among those notions, the concept of weak observability relaxes classical observability in the sense of requiring invertibility of the output map only in a neighbourhood of a given state. One of the strengths of this notion is that it is particularly amenable to algebraic (rank) conditions which generalize the Kalman rank conditions for (6) available in the linear case. Let us briefly recall the discussions from [45, 60] with slight adjustments to account for the simplified setup considered here and define the Lie derivative of a scalar (component) function $h_{i}\colon\mathcal{M}\to\mathbb{R},i=1,\dots,m$ along the uncontrolled vector field $f\colon\mathcal{M}\to\mathcal{M}$ as
$L_{f}h_{i}(x)=\nabla h_{i}(x)^{\top}f(x).$ (26)
@@
$\mathrm{d}\mathcal{G}(x):=\{z\in\mathbb{R}^{n}\ |\ z=\nabla\phi(x),\phi\in\mathcal{G}\},$ (27)
we have the following well-known results, see [45, Theorem 3.1, Theorem 3.12].
@@
Unfortunately, the above rank conditions are not easily verifiable in all but the simplest examples. Moreover, the conditions only imply weak observability and therefore generally do not ensure invertibility of the output map, see [45, Example 3.3]. Even in the case of an invertible output map, explicit reconstruction formulas in terms of (nonlinear) observability Gramians do not exist in the general nonlinear case. On the other hand, it is known that observability is a generic property that is guaranteed to hold almost surely if arbitrarily smooth vector fields are considered [61]. The latter reference further shows that observability can be achieved in a (practically relevant) setup where only finitely many output samples rather than a continuous output trajectory are available. In particular, it is shown that the number of samples has to be at least $12n+$ 1 for an $n$-dimensional state space. Finally, let us also point to [62] which considers potential extensions of linear observability measures in terms of local generalizations of the observability Gramian and related cost functionals.
A.3 Autoregressive Training with GANs
@@
$=\mathbb{E}_{x_{0}\sim\mu,t\sim U(\Gamma_{T}\setminus(\Gamma_{\Delta(k-1)}\cup\{T\}))}\left[\sup_{v\in\mathbb{R}^{m},||v||=1}v\cdot\left(\eta_{g}(y_{\mathrm{seq}}(t;x_{0}))-y(t+\Delta;x_{0})\right)\right]$
@@
where the last step is possible as $v_{g}=\texttt{norm}(\eta_{g}(y_{\text{seq}})-g(y_{\text{seq}}))$ is an explicit point wise maximizer, where $\0\texttt{norm}(v)=\frac{v}{||v||},v\in\mathbb{R}^{m}\textbackslash\{0\}$. Note that here again we used existence of $g$ and hence observability of the dynamic system. The idea of adversarial training is to learn the function $v_{g}$ with a neural network from a sufficiently large hypothesis space $\mathcal{H}^{v}_{g,v,T}$ of neural networks with norm activation in the last layer. Consider the min-max optimisation problem
@@
^$\hat{\mathcal{L}}_{g,N,T}^{v}(\eta_{g},\eta_{g}^{v})=\frac{1}{N\left(\dfrac{T}{\Delta}-k\right)}\sum^{N}_{j=1}\sum_{t\in\Gamma_{T}\setminus(\Gamma_{k\Delta}\cup\{T\}))}[\eta^{v}_{g}(y_{\text{seq}}(t;x^{(j)}_{0}))\cdot\eta_{g}(y_{\text{seq}}(t;x^{(j)}_{0}))$
@@
Given ^$\hat{\mathcal{L}}_{g}^{v}$, the training of the ‘critic’ network $\hat{\eta}_{g}^{v}$ and the generator $\hat{\eta}_{g}$ is done in the usual adversarial fashion
@@
For $\eta_{G}$ we can set up an analogous adversarial training scheme. Using the decomposition for adversarial learning [63, 64, 65] an adversarial treatment of PAC-learning is feasible as well. This however goes beyond the scope of this article. We also note that in the LongVideoGAN [48] experiments, adversarial training is performed with respect to a ’discriminator’ instead of a ’censor’ network , see [66] for the difference.
A.4 Discretization of the PDE
In this section we aim to discretise the PDEs (Eqs. 11 and 13) and give an account of our numerical solvers.
A.4.1 Linear Equations
@@
The left and right discrete derivatives are (in the $x$-direction) [67]
$D_{x}^{+}u_{i,j}:=\frac{u_{i+1,j}-u_{i,j}}{\Delta x},\;\;\;D_{x}^{-}u_{i,j}:=\frac{u_{i,j}-u_{i-1,j}}{\Delta x}$ (28)
@@
$(\nabla\cdot\left(a\nabla\right)u)_{i}$ $=\frac{\partial}{\partial x}\left(a\frac{\partial}{\partial x}\right)u+\frac{\partial}{\partial y}\left(a\frac{\partial}{\partial y}\right)u$ (29)
$\approx\sum^{n^{2}}_{j=1}\left[\sum_{k=1}^{n^{2}}\left(D_{x}^{-}\right)_{ik}a_{k}\left(D_{x}^{+}\right)_{kj}+\sum_{k=1}^{n^{2}}\left(D_{y}^{-}\right)_{ik}a_{k}\left(D_{y}^{+}\right)_{kj}\right]u_{j}$ (30)
@@
For an appropriately generated positive definite vector $a$; see A.6. In the case $a$ is a vector of ones, this expression reduces to the standard five-point stencil. We write $u_{i,j}$ which lives in the torus defined on $[-1,1]^{2}$ as a vector $u_{k}$, with dimension $n^{2}$, and as such a mapping (with appropriately wrapping indices) $(i,j)\mapsto(ni+j)$ is required to perform the matrix multiplications defined in 31.
@@
$u_{i,j}^{\tau+1}=\Delta t\left(A_{h}u_{i,j}^{\tau}\right)+u_{i,j}^{\tau}$ (32)
@@
The discretisation for the wave equation is similar, but instead requires the first order formulation 12, which gives the update rule for the full state space $u^{\prime}_{i,j}=(u,v)_{i,j}$ as
@@
A.4.2 Nonlinear Equations
The KS dataset was generated by numerically integrating the governing partial differential equation 13 using a fourth-order Exponential Time Differencing Runge-Kutta method (ETDRK4), as described in [50]. This method is particularly well-suited for stiff PDEs such as the KS equation due to its ability to handle the linear stiff components analytically while treating the nonlinear terms explicitly. The spatial discretization was performed using Fourier spectral methods under periodic boundary conditions. The scheme uses a precomputed set of coefficients derived from contour integrals in the complex plane, which helps resolves some numerical instability issues.
@@ A.4.2 Nonlinear Equations
A.5 Local Observability of the KS Equation
@@
The KSE (13) in one spatial dimension $\frac{\partial u}{\partial t}+u\,\frac{\partial u}{\partial x}+\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{4}u}{\partial x^{4}}=0$ with periodic boundary conditions is solved via the ETDRK2-Method [68]. We define a tokenized state $y(t)\in\mathbb{R}^{m}$ by averaging $u(x,t)$ over sliding spatial windows with stride $p=5$. After an initial buffer of snapshots sufficient to approximate Lie derivatives up to order $p=5$ via forward finite differences, the observability matrix $\mathcal{O}(t_{k})$ is constructed by stacking these Lie-derivative approximations.
This yields a local linear approximation of the observability matrix $\mathcal{O}(x_{t})$ at $x_{t}$ whose sign-preserving $\log\det(\mathcal{O}(x_{t}))$ remains finite, indicating full rank of the observability matrix $\mathcal{O}(x_{t})$ over the time interval and hence local weak observability of the dynamical system after a burn-in period which can be interpreted as the time of transition from an initially almost constant state to the onset of the chaotic dynamics.
Figure 7: Time trace of $\log(\det\mathcal{O}(x_{t}))$ and its centered 50-step rolling average. Here $L=80$, $u(x,0)=\sin(7\pi x/L)$; spectral discretization with $N=200$ modes and ETDRK2 time stepping ($\Delta t=0.01$, $T=10000$ steps); coarse state $y\in\mathbb{R}^{N/p}$ via averaging over sliding spatial windows of size $p=5$ ($N/p=40$); first $p=5$ time-derivatives by forward differences. The bounded $\log\det$ indicates $\mathrm{rank}\,\mathcal{O}(x_{t})=N$.
A.6 Further Dataset Details

### v2

@@ Appendix A Extraneous Proofs
In the case that (1) defines an ergodic dynamical system with invariant measure $\mu$, we can replace the large sample limit in the number of trajectories $N(\varepsilon,\delta)\to\infty$ as $\varepsilon,\delta\downarrow 0$ with the large time limit $T(\varepsilon,\delta)\to\infty$ keeping $N=1$ fixed, see drygala2022generative for this approach in a related situation based on Birkhoff’s ergodic theorem.
Appendix B Continuous Time Control Theory
In continuous time control theory, Definition 1 (i) is changed by demanding $y(t;x_{0})=y(t;x_{0})$ for $t\in[0,T]$ instead of only for $t\in\Gamma_{T}$. Note that the dynamical system (1) is a special case of Controllability_and_Observability since the distinguishing control is assumed to be the trivial one. Let us emphasize that for nonlinear systems this generally diminishes otherwise existent observability properties, cf. (Cas82, Section 5). On the other hand, considering uncontrolled systems has the advantage of characterizing observability equivalently in terms of the injectivity of the so-called output map
$H\colon\mathcal{X}\to L^{2}(0,T;\mathbb{R}^{m}),\ \ H(x_{0}):=h(x(\cdot;x_{0}))=y(\cdot;x_{0})$
which assigns to a given initial state its corresponding output trajectory. As a consequence, observability of $(f,h)$ implies the existence of an inverse $H^{-1}\colon H(\mathcal{X})\to\mathcal{X}$. Composing this function with the flow of (1) moreover allows us to define a reconstruction map
$G\colon H(\mathcal{X})\to L^{2}(0,\tilde{T};\mathbb{R}^{n}),\ \ G(y):=(S_{\tilde{T}}\circ H^{-1})(y)=x(\cdot;H^{-1}(y))$
@@ Appendix B Continuous Time Control Theory
$H^{-1}y=Q(\tilde{T})^{-1}\int_{0}^{\tilde{T}}e^{A^{\top}t}h^{\top}y(t)\,\mathrm{d}t,\quad Gy=e^{A\cdot}Q(\tilde{T})^{-1}\int_{0}^{\tilde{T}}e^{A^{\top}t}h^{\top}y(t)\,\mathrm{d}t$
where the positive definiteness of the observability Gramian $Q(\tilde{T}):=\int_{0}^{\tilde{T}}e^{A^{\top}s}h^{\top}he^{As}\,\mathrm{d}s$ is ensured by the observability of the system, see (Son98, Section 6.3).
In Controllability_and_Observability, other notions of observability have been discussed. Among those notions, the concept of weak observability relaxes classical observability in the sense of requiring invertibility of the output map only in a neighbourhood of a given state. One of the strengths of this notion is that it is particularly amenable to algebraic (rank) conditions which generalize the Kalman rank conditions for (6) available in the linear case. Let us briefly recall the discussions from Controllability_and_Observability; Cas82 with slight adjustments to account for the simplified setup considered here and define the Lie derivative of a scalar (component) function $h_{i}\colon\mathcal{M}\to\mathbb{R},i=1,\dots,m$ along the uncontrolled vector field $f\colon\mathcal{M}\to\mathcal{M}$ as
$L_{f}h_{i}(x)=\nabla h_{i}(x)^{\top}f(x).$
@@
$\mathrm{d}\mathcal{G}(x):=\{z\in\mathbb{R}^{n}\ |\ z=\nabla\phi(x),\phi\in\mathcal{G}\},$
we have the following well-known results, see (Controllability_and_Observability, Theorem 3.1, Theorem 3.12).
@@
Unfortunately, the above rank conditions are not easily verifiable in all but the simplest examples. Moreover, the conditions only imply weak observability and therefore generally do not ensure invertibility of the output map, see (Controllability_and_Observability, Example 3.3). Even in the case of an invertible output map, explicit reconstruction formulas in terms of (nonlinear) observability Gramians do not exist in the general nonlinear case. On the other hand, it is known that observability is a generic property that is guaranteed to hold almost surely if arbitrarily smooth vector fields are considered (Aey81). The latter reference further shows that observability can be achieved in a (practically relevant) setup where only finitely many output samples rather than a continuous output trajectory are available. In particular, it is shown that the number of samples has to be at least $2n+1$ for an $n$-dimensional state space. Finally, we also point to Zen18 which considers potential extensions of linear observability measures in terms of local generalisations of the observability Gramian and related cost functionals.
Appendix C Comparison to Other Methods
Ours (full model) DeepONet FNO
KSE (200) $\mathbf{5.7109}$ $6.1465$ $6.6150$
Table 2: Results on the KSE dataset for the time correlation metric. The Euclidean distance between each point of the time correlation curves of the simulations and the model is measured and summed, up to 200 frames.
We include a table below to compare the performance of the models on OOD initial conditions. We tested on three OOD initial conditions:
  1. 1.
    The field $z=(x(1-x)y(y-1))^{2},\;\;\;0<x,y<1$, with $z$ normalised to [1,1][-1,1], labelled as ‘eq’ in the table. This gives a smooth, non-trivial initial condition with zero values at the boundaries of the domain;
  2. 2.
    A single central point source ‘point’, where the centre of the grid is set to $1$ and everywhere else left as 0. This tests a highly discontinuous initial condition;
  3. 3.
    A sin wave ‘sin’, given by $z=\sin{(2\pi x)}\sin{(2\pi y)},\;\;\;0<x,y<1$
All three conditions are evaluated on a meshgrid, discretised to the resolution required by the dataset. The initial conditions are evolved using the same numerical techniques given in Section E.
Dataset Condition Frame DeepONet FNO Ours
L1 MSE L1 MSE L1 MSE
Heat eq 1 0.01217 0.0002134 0.005122 $\mathbf{5.544\times 10^{-5}}$ 0.1145 0.02102
Heat eq 200 7.133 1397 $7.175\times 10^{5}$ $4.264\times 10^{12}$ 0.2148 0.08548
Heat eq 1000 Diverged Diverged Diverged Diverged 0.669 0.7396
Heat point 1 0.00401 $\mathbf{2.559\times 10^{-5}}$ 0.005657 $6.812\times 10^{-5}$ 0.04133 0.009935
Heat point 200 10.74 2707 0.2952 0.1134 0.4876 0.4414
Heat point 1000 Diverged Diverged Diverged Diverged 1.525 3.84
Heat sin 1 0.007225 0.0001037 0.004157 $\mathbf{2.608\times 10^{-5}}$ 0.1241 0.02409
Heat sin 200 $4.219\times 10^{5}$ $8.944\times 10^{11}$ $1.6\times 10^{7}$ $1.242\times 10^{15}$ 0.03128 0.001517
Heat sin 1000 Diverged Diverged Diverged Diverged 0.01401 0.0003164
Wave eq 1 0.02106 0.0004665 0.009501 0.0001064 0.04258 0.003717
Wave eq 200 0.7729 0.6937 $4.55\times 10^{7}$ $1.4\times 10^{16}$ 1.4 4.243
Wave eq 1000 Diverged Diverged Diverged Diverged 6.786 64.6
Wave point 1 0.003903 $2.37\times 10^{-5}$ 0.001271 $\mathbf{3.376\times 10^{-6}}$ 0.0446 0.008845
Wave point 200 $5.508\times 10^{6}$ $2.211\times 10^{15}$ $7.155\times 10^{6}$ $3.807\times 10^{14}$ 1.026 17.74
Wave point 1000 Diverged Diverged Diverged Diverged 13.81 3353
Wave sin 1 0.007625 $8.853\times 10^{-5}$ 0.004253 $\mathbf{3.937\times 10^{-5}}$ 0.04471 0.003131
Wave sin 200 $6.538\times 10^{5}$ $2.076\times 10^{13}$ $3.475\times 10^{6}$ $1.028\times 10^{14}$ 0.8098 1
Wave sin 1000 Diverged Diverged Diverged Diverged 3.57 19.5
KSE eq 1 0.0111 0.000147 0.02004 0.0005884 0.2302 0.1113
KSE eq 200 2.516 8.534 0.02438 0.0008629 0.3907 0.2384
KSE point 1 0.01244 0.0001977 0.07488 0.01854 0.1859 0.1057
KSE point 200 2.584 10.24 2.197 8.133 0.3008 0.1318
KSE sin 1 0.0111 0.000147 0.02006 0.0005886 0.1696 0.1217
KSE sin 200 2.514 8.527 0.02792 0.001137 0.2992 0.132
Table 3: Performance comparison between DeepONet, FNO, and our model across datasets, initial conditions, and generation horizon (frame). Best values in each row are highlighted in bold.
Our model is the most consistent, but is occasionally beaten on the 1-frame ahead metrics. However, as the models roll out to further time horizons, our model is generally the best. As with the in-distribution metrics, none of the alternative models survive to 1000 frames.
Appendix D Autoregressive Training with GANs
@@
[…]
@@
[…]
@@
$\hat{\mathcal{L}}_{g,N,T}^{v}(\eta_{g},\eta_{g}^{v})=\frac{1}{N\left(\dfrac{T}{\Delta}-k\right)}\sum^{N}_{j=1}\sum_{t\in\Gamma_{T}\setminus(\Gamma_{k\Delta}\cup\{T\}))}[\eta^{v}_{g}(y_{\text{seq}}(t;x^{(j)}_{0}))\cdot\eta_{g}(y_{\text{seq}}(t;x^{(j)}_{0}))$
@@
Given $\hat{\mathcal{L}}_{g}^{v}$, the training of the ‘critic’ network $\hat{\eta}_{g}^{v}$ and the generator $\hat{\eta}_{g}$ is done in the usual adversarial fashion
@@
For $\eta_{G}$ we can set up an analogous adversarial training scheme. Using the decomposition for adversarial learning (asatryan2020convenient; biau2021some; drygala2025learning) an adversarial treatment of PAC-learning is feasible as well. This however goes beyond the scope of this article. We also note that in the LongVideoGAN (Long_video_gan) experiments, adversarial training is performed with respect to a ‘discriminator’ instead of a ‘censor’ network, see arjovsky2017wasserstein for the difference.
Appendix E Discretization of the PDE
In this section we aim to discretise the PDEs (Eqs. 10 and 12) and give an account of our numerical solvers.
E.1 Linear Equations
@@
The left and right discrete derivatives are (in the $x$-direction) (finitedifferencemethods)
$D_{x}^{+}u_{i,j}:=\frac{u_{i+1,j}-u_{i,j}}{\Delta x},\;\;\;D_{x}^{-}u_{i,j}:=\frac{u_{i,j}-u_{i-1,j}}{\Delta x}$
@@
$(\nabla\cdot\left(a\nabla\right)u)_{i}$ $=\frac{\partial}{\partial x}\left(a\frac{\partial}{\partial x}\right)u+\frac{\partial}{\partial y}\left(a\frac{\partial}{\partial y}\right)u$
$\approx\sum^{n^{2}}_{j=1}\left[\sum_{k=1}^{n^{2}}\left(D_{x}^{-}\right)_{ik}a_{k}\left(D_{x}^{+}\right)_{kj}+\sum_{k=1}^{n^{2}}\left(D_{y}^{-}\right)_{ik}a_{k}\left(D_{y}^{+}\right)_{kj}\right]u_{j}$
@@
For an appropriately generated positive definite vector $a$; see Section G. In the case $a$ is a vector of ones, this expression reduces to the standard five-point stencil. We write $u_{i,j}$ which lives in the torus defined on $1,1]^{2}$ as a vector $u_{k}$, with dimension $n^{2}$, and as such a mapping (with appropriately wrapping indices) $(i,j)\mapsto(ni+j)$ is required to perform the matrix multiplications defined in Eq. 22.
@@
$u_{i,j}^{\tau+1}=\Delta t\left(A_{h}u_{i,j}^{\tau}\right)+u_{i,j}^{\tau}$
@@
The discretisation for the wave equation is similar, but instead requires the first order formulation 11, which gives the update rule for the full state space $u^{\prime}_{i,j}=(u,v)_{i,j}$ as
@@
E.2 Nonlinear Equations
The KS dataset was generated by numerically integrating the governing partial differential equation 12 using a fourth-order Exponential Time Differencing Runge-Kutta method (ETDRK4), as described in etdrk4. This method is particularly well-suited for stiff PDEs such as the KS equation due to its ability to handle the linear stiff components analytically while treating the nonlinear terms explicitly. The spatial discretization was performed using Fourier spectral methods under periodic boundary conditions. The scheme uses a precomputed set of coefficients derived from contour integrals in the complex plane, which helps resolves some numerical instability issues.
@@ E.2 Nonlinear Equations
Appendix F Local Observability of the KS Equation
@@
The KSE (12) in one spatial dimension $\frac{\partial u}{\partial t}+u\,\frac{\partial u}{\partial x}+\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{4}u}{\partial x^{4}}=0$ with periodic boundary conditions is solved via the ETDRK2-Method (ETDRK2). We define a tokenized state $y(t)\in\mathbb{R}^{m}$ by averaging $u(x,t)$ over sliding spatial windows with stride $p=5$. After an initial buffer of snapshots sufficient to approximate Lie derivatives up to order $p=5$ via forward finite differences, the observability matrix $\mathcal{O}(t_{k})$ is constructed by stacking these Lie-derivative approximations.
This yields a local linear approximation of the observability matrix $\mathcal{O}(x_{t})$ at $x_{t}$ whose sign‐preserving $\log\det(\mathcal{O}(x_{t}))$ remains finite, indicating full rank of the observability matrix $\mathcal{O}(x_{t})$ over the time interval and hence local weak observability of the dynamical system after a burn-in period which can be interpreted as the time of transition from an initially almost constant state to the onset of the chaotic dynamics.
Figure 7: Time trace of $\log(\det\mathcal{O}(x_{t}))$ and its centered 50‐step rolling average. Here $L=80$, $u(x,0)=\sin(7\pi x/L)$; spectral discretization with $N=200$ modes and ETDRK2 time stepping ($\Delta t=0.01$, $T=10000$ steps); coarse state $y\in\mathbb{R}^{N/p}$ via averaging over sliding spatial windows of size $p=5$ ($N/p=40$); first $p=5$ time‐derivatives by forward differences. The bounded $\log\det$ indicates $\mathrm{rank}\,\mathcal{O}(x_{t})=N$.
Appendix G Further Dataset Details